Please help. Also so how the problem was solved.
Tickets for a concert were sold to adults for $3 and to students for $2. If The total receipts were $824 and twice as many adults tickets as students tickets were sold, then how many of each were sold?

Answer by checkley77(12569)

2x*3+x*2=824
6x+2x=824
8x=824
x=824/8
x=103 student tickets were sold
28103=206 adult tickets were sold.
Proof:
2*103*3+103*2=824
618+206=824
824=824